# DOMAIN: milDly cOnservative Model-bAsed offlINe reinforcement learning

### v3

@@ Abstract
Model-based reinforcement learning (RL), which learns environment model from offline dataset and generates more out-of-distribution model data, has become an effective approach to the problem of distribution shift in offline RL. Due to the gap between the learned and actual environment, conservatism should be incorporated into the algorithm to balance accurate offline data and imprecise model data. The conservatism of current algorithms mostly relies on model uncertainty estimation. However, uncertainty estimation is unreliable and leads to poor performance in certain scenarios, and the previous methods ignore differences between the model data, which brings great conservatism. Therefore, this paper proposes a milDly cOnservative Model-bAsed offlINe RL algorithm (DOMAIN) without estimating model uncertainty to address the above issues. DOMAIN introduces adaptive sampling distribution of model samples, which can adaptively adjust the model data penalty. In this paper, we theoretically demonstrate that the Q value learned by the DOMAIN outside the region is a lower bound of the true Q value, the DOMAIN is less conservative than previous model-based offline RL algorithms and has the guarantee of safety policy improvement. The results of extensive experiments show that DOMAIN outperforms prior RL algorithms on the D4RL dataset benchmark.
@@ I Introduction
Reinforcement learning (RL) has been widely applied in various fields, including robot control [1, 2], autonomous driving [3], recommendation systems [4], and resource scheduling [5]. The paradigm of online RL involves an agent interacting with environment directly through trial-and-error[6]. However, online RL suffers from low sampling efficiency and high interaction cost, making it impractical in certain scenarios [7]. Offline RL can effectively address the above problems by learning policy from offline dataset without interaction with the environment.
Offline RL faces a significant challenge due to distribution shift, which is caused by difference between the policy in offline dataset and those learned by the agent. The selection of out-of-distribution (OOD) data introduces extrapolation error that may lead to unstable training [9]. To address the above issue, two types of methods have been developed: policy constraints [10, 11, 12] and value penalization [13, 14]. The goal of the above methods is to minimize the discrepancy between the learned policy and that in the dataset. While they achieve remarkable performance, these methods are model-free and fail to consider the effects of OOD data, hampering the exploration and performance improvement.
Model-based offline RL, which enables agent to continually interact with a trained environment model, can broaden data coverage and enhance exploration in OOD region [16]. However, due to the gap between the estimated and true environment model, the generated data maybe inaccurate, that is the generated model data is inaccurate, which brings risks for agent training. The current methods often perform conservative policy optimization to solve this problem. The typical method is estimating model uncertainty [19, 20, 39] or penalizing state-action pairs with high uncertainty [24, 25]. However, estimating model uncertainty through neural networks brings prediction errors and low reliability. To address this problem, Yu et al. [21] proposed the COMBO algorithm based on CQL [13], which regularizes model data to suppress unreliable data and enhance reliable offline data.
@@
However, when regularizing model data, the current model-based offline RL methods fail to consider the differences between the model data and are overly conservative. For example, for precise model data, penalizing them as if they were imprecise model data can lead to underutilization of the precise data, thereby increasing the conservativeness of the algorithm. Therefore, we design an adaptive sampling distribution of model data, and propose a mildly conservative model-based offline RL algorithm (DOMAIN) without estimating model uncertainty (as illustrated in Fig. 1). Here, the mildly conservatism is reflected as: higher exploration in OOD region and better performance of agent. This algorithm imposes heavier penalties on model data with high errors while reducing the penalty for accurate model data. Our main contributions are summarized as follows:
@@
    Theoretical analysis demonstrates that the value function learned by the DOMAIN in the OOD region is a lower bound of the true value function, and DOMAIN has a safety policy improvement guarantee.
@@
    The results of extensive experiment indicate that DOMAIN can outperform previous RL algorithms on the D4RL dataset benchmark.
@@
Notably, the exploration risk for unknown region can be reduced through underestimating the value of inaccurate model data. Through reducing the penalty on the $Q$ value of model data with small error, the agent can choose action $a$ with higher reward $r$ by $\arg\max_{a}Q(s,a)$, thus improving performance of the learned policy. The framework of this paper is as follows: Section 2 introduces the fundamental theory of RL. Section 3 shows the framework and implementation details of the DOMAIN algorithm. Section 4 gives some relevant theoretical analysis. Section 5 presents the algorithm performance on D4RL tasks. Finally, Section 6 summarizes the entire work.
II Preliminaries and related works
II-A Reinforcement Learning
@@
The framework of RL is based on the Markov Decision Process (MDP) $\mathcal{M}=\left(\mathcal{S},\mathcal{A},r,T_{\mathcal{M}},\rho_{0},\gamma\right)$, where $\mathcal{S}$ is the set of states, $\mathcal{A}$ is the set of actions, $r(s,a)$ is the reward function, $T_{\mathcal{M}}\left(s^{\prime}\mid s,a\right)$ represents the dynamic system under $\mathcal{M}$, $\rho_{0}$ is the initial state distribution, and $\gamma\in(0,1)$ is the discount factor [8]. The goal of RL is to learn a policy that maximizes the cumulative reward $J(\mathcal{M},\pi)=\mathbb{E}_{(s,a)\sim d_{\mathcal{M}}^{\pi}(s,a)}[r(s,a)]/1-\gamma$, where $d_{\mathcal{M}}^{\pi}(s,a)=d_{\mathcal{M}}^{\pi}(s)\pi(a\mid s)$ is the marginal distribution of states and actions under the learned policy $\pi$, and $d_{\mathcal{M}}^{\pi}(s)$ is the discounted marginal state distribution under the learned policy $\pi$.
@@
II-B Offline RL
The goal of offline RL is to learn a policy $\pi_{\theta}(\cdot\mid s)$ from an offline dataset $\mathcal{D}=\left\{\left(s_{i},a_{i},r_{i},s_{i}^{\prime}\right)\right\}_{i=1}^{n}$ without interacting with the environment, to maximize the expected discounted reward. However, there exists the distribution shift between the learned policy $\pi_{\theta}(\cdot\mid s)$ and the behavioral policy $\pi_{b}(\cdot\mid s)$ of the dataset, leading to unstable training of the agent [8]. To alleviate this issue, previous studies [10, 11, 12, 13, 14] have introduced regularization on the value function of OOD actions to reduce the probability of selecting OOD actions. Here primarily focuses on CQL[13]. The formulas for policy improvement and policy evaluation are as follows:
@@ II-B Offline RL
II-C Model-based Offline RL
Model-based offline RL refers to training an environment model $\widehat{T}_{\phi}\left(s^{\prime},r\mid s,a\right)$ using the offline dataset, where the model is typically parameterized as Gaussian distribution $\widehat{T}_{\phi}(s^{\prime},r\mid s,a)=\mathcal{N}[\mu_{\phi}(s,a),\Sigma_{\phi}(s,a)]$. The loss function for training the environment model employs maximum likelihood estimation as: $\mathcal{L}_{\mathcal{M}}(\phi)=-\mathbb{E}_{\left(s,a,r,s^{\prime}\right)\sim\mathcal{D}}[\log\widehat{T}_{\phi}(s^{\prime},r\mid s,a)]$. With this learned model, an estimated MDP $\widehat{\mathcal{M}}$ can be obtained. Subsequently, the agent interacts with the learned policy $\widehat{\pi}$ and the environment model $\widehat{T}_{\phi}$ to generate model data $\mathcal{D}_{\text{model }}$ through $H$-step rollouts. The policy learning is then performed using the combined dataset $\mathcal{D}_{\text{offline }}\cup\mathcal{D}_{\text{model }}$, where the data is sampled from model data $\mathcal{D}_{\text{model }}$ with the probability $f,f\in[0,1]$, and from the offline data $\mathcal{D}_{\text{offline }}$ with the probability $1-f$.
However, offline dataset fails to cover the entire state-action space, resulting in disparities between the learned environment model $\widehat{T}_{\phi}\left(s^{\prime},r\mid s,a\right)$ and the actual environment model $T\left(s^{\prime},r\mid s,a\right)$. To address this challenge, Janner et al. [23] employed a model ensemble approach to reduce biases introduced by probabilistic networks. Yu et al. [19] proposed the MOPO algorithm, which used model uncertainty as reward penalizes. Yu et al. [21] considered the estimation errors of neural networks for model uncertainty and utilized value function regularization, aiming to push down unreliable model data and push up reliable offline data. However, the above methods fail to consider the difference between the model data and overly conservative, thus the algorithm DOMAIN is proposed in this paper.
III Methodology
This section gives the details of mildly conservative model-based offline RL (DOMAIN) algorithm, including mildly conservative estimation for value function and adaptive value regularization on model data.
III-A Mildly Conservative Value Estimation
Conservative policy evaluation within the framework of DOMAIN algorithm can be described as follows: Given an offline dataset $\mathcal{D}_{\text{offline }}=\left\{\left(s_{i},a_{i},r_{i},s_{i}^{\prime}\right)\right\}_{i=1}^{n}$, a learned policy $\widehat{\pi}$, and a learned environmental model $\widehat{T}_{\phi}$, the primary goal is to obtain a conservative estimation of $Q(s,a)$. To accomplish this, the policy evaluation step in the DOMAIN employs the CQL framework (see Eq. (2)). This principle can be mathematically expressed as:
@@ III-A Mildly Conservative Value Estimation
where $\widehat{\mathcal{T}}^{\pi}$ represents the actual Bellman operator used in the algorithm, while $\mathcal{T}_{\overline{\mathcal{M}}}^{\pi}$ and $\mathcal{T}_{\widehat{\mathcal{M}}}^{\pi}$ represent the operator applied under the empirical MDP and the learned model, respectively.
The value function $Q$ and the policy $\pi$ in Eq. (4) can be approximated using neural networks, denoted as $Q_{\omega}$ and $\pi_{\theta}$, respectively. The value function update follows the SAC framework [27], and thus the Bellman operator can be further expressed as:
@@
where $\alpha$ is the entropy regularization coefficient, determining the relative importance of entropy. By employing a mildly conservative policy evaluation, the policy improvement scheme can be obtained:
@@
where $Z$ is the normalization factor. Substituting the solution obtained from solving Eq. (7) back into Eq. (3), the following expression can be obtained:
@@
III-B Adaptive Value Regularization on Model Data
When performing policy evaluation, we want to assign higher penalty coefficients to larger error model data, thereby yielding an increased value for $\omega(s,a)$. Drawing inspiration from [28, 29], the KL divergence is employed to quantify the extent of model data errors:

### v4

@@ Abstract
Model-based reinforcement learning (RL), which learns an environment model from the offline dataset and generates more out-of-distribution model data, has become an effective approach to the problem of distribution shift in offline RL. Due to the gap between the learned and actual environment, conservatism should be incorporated into the algorithm to balance accurate offline data and imprecise model data. The conservatism of current algorithms mostly relies on model uncertainty estimation. However, uncertainty estimation is unreliable and leads to poor performance in certain scenarios, and the previous methods ignore differences between the model data, which brings great conservatism. To address the above issues, this paper proposes a milDly cOnservative Model-bAsed offlINe RL algorithm (DOMAIN) without estimating model uncertainty, and designs the adaptive sampling distribution of model samples, which can adaptively adjust the model data penalty. In this paper, we theoretically demonstrate that the Q value learned by the DOMAIN outside the region is a lower bound of the true Q value, the DOMAIN is less conservative than previous model-based offline RL algorithms, and has the guarantee of safety policy improvement. The results of extensive experiments show that DOMAIN outperforms prior RL algorithms and the average performance has improved by $1.8\%$ on the D4RL benchmark.
@@ I Introduction
Reinforcement learning (RL) has been widely applied in various fields, including robot manipulation control [1, 2], medical robotics control [3, 4], underwater robotics control [5, 6], and resource scheduling [7]. RL learns the optimal policy through the agent interacting with the environment constantly, that is, trial-and-error [8]. However, agents directly interacting with the environment may bring low sampling efficiency and high interaction cost [9], which makes it impractical in certain real-world scenarios. Therefore, offline RL has been widely studied in recent years.
Offline RL, which learns the policy froma static dataset without interaction with the environment, has become an effective way to address the above problem. However, since limited coverage of the collected dataset, offline RL suffers from distribution shift, which is caused by the difference between the policy in the offline dataset and those learned by the agent [11]. The selection of out-of-distribution (OOD) data introduces extrapolation error, which may lead to unstable training and degrade agent performance [16]. Offline RL can be divided into two parts: model-free RL and model-based RL, which take different approaches to solve the above problems.
Model-free offline RL, independent of environment model, addresses the distribution shift challenge through two principal techniques: policy constraints [17, 18, 19] and value penalization [20, 21]. The goal of the above methods is to minimize the discrepancy between the learned policy and that in the offline dataset. While they achieve remarkable performance, these methods fail to consider the effects of OOD data, hampering the exploration and performance improvement. Lyu et al. [22] proved that the performance can be effectively improved through exploring the OOD region. Therefore, some data augmentation techniques are studied to improve performance.
Model-based offline RL, which trains an environment model through the offline dataset, enables the agent to continually interacting with the trained environment model to broaden data coverage and enhance exploration in the OOD region [23]. The expansion of OOD region data can effectively mitigate distribution shift. However, due to the gap between the estimated and true environment model, the generated data may be inaccurate, that is, the generated model data is inaccurate, which brings risks for agent training. The current methods solve this problem by performing conservative policy optimization. There are two types of methods to achieve conservatism. The first type involves imposing reward penalties on OOD data, either directly through uncertainty quantification [26, 27, 31] or indirectly via value function regularization [28]. The second type adopts a robust adversarial training framework, which enforces conservatism by training an adversarial environment model [36, 37].
The uncertainty quantification of the first type [26, 31], estimated by neural networks, suffers from low reliability and reduces agent performance. Therefore, Yu et al. [28] proposed the COMBO algorithm based on CQL [20], that is, the value regularization of the first type, which regularizes model data to suppress unreliable data and enhance reliable offline data without estimating uncertainty. The second type of methods [36, 37] require continuous updates to the environment model during policy learning, leading to increased computational demands. Table I summarizes the recent model-based offline RL methods by the type of algorithm and computation cost. In this paper, we focus on the value regularization method of the first type.
@@
TABLE I: The summary of recent model-based offline RL research.
Methods Uncertainty Regularization Adversarial Model data Computation
qualification indirectly training difference complexity
MOPO [26] low
MOReL[27] low
MAPLE[31] low
RAMBO[36] high
ARMOR[37] high
COMBO[28] low
DOMAIN low
For model-based offline RL, when regularizing model data to perform conservative policy, the current methods fail to consider the differences between the model data, that is, the errors ofthe model data are different, which brings overly conservative policy. For example, for precise model data, penalizing them as if they were imprecise model data can lead to underutilization of the precise data, thereby increasing the conservativeness of the algorithm and reducing agent performance. Motivated by this problem, we design an adaptive sampling distribution of model data, where the regularization for model data can be adjusted adaptively. Then, a mildly conservative model-based offline RL algorithm (DOMAIN) is proposed without estimating model uncertainty (as illustrated in Fig. 1). Our main contributions are summarized as follows:
@@
    Theoretical analysis demonstrates that the value function learned by the DOMAIN in the OOD region is a lower bound of the true value function, and the DOMAIN has a safety policy improvement guarantee.
@@
    The results of an extensive experiment indicate that DOMAIN can outperform previous RL algorithms on the D4RL benchmark.
@@
Here, the mild conservatism of DOMAIN is reflected as: higher exploration in the OOD region and better performance of the agent. This algorithm imposes heavier penalties on model data with high errors while reducing the penalty for accurate model data. Notably, the exploration risk for the unknown region can be reduced through underestimating the value of inaccurate model data. Through reducing the penalty on the $Q$ value of model data with small error, the agent can choose action $a$ with higher reward $r$ by $\arg\max_{a}Q(s,a)$, thus improving the performance of the learned policy.
The framework of this paper is as follows: Section II provides the detailed literature for model-free and model-based offline RL. Section III introduces the fundamental theory of RL. Section IV shows the framework and implementation details of the DOMAIN algorithm. Section V gives some relevant theoretical analysis for DOMAIN. Section VI presents the algorithm performance of DOMAIN on D4RL tasks. Finally, Section VII summarizes the entire work.
II Related works
This section presents a brief overview for the model-free offline RL and model-based offline RL algorithms, and the differences and advantages between our work and the previous related works.
Model-free Offline RL. The current model-free offline RL algorithms take different constraints or regularization techniques to mitigate the extrapolation error. Huang et al. [29] proposed a mild policy evaluation method by constraining the difference between the $Q$ values of actions supported by the target policy and those of actions contained within the offline dataset. Yang et al. [30] treated different samples with different policy constraint intensities, which effectively improves the agent’s performance. Huang et al. [12] proposed the de-pessimism operator to estimate $Q$ values through judging whether the actions are in the OOD region. Cao et al. [13] simultaneously performed behavior regularization on policy evaluation and policy improvement to improve performance. However, the main challenge in offline RL arises from the limited data coverage. Model-free offline RL algorithms rely solely on offline datasets to learn policies, which limits the agent’s ability to explore. Therefore, some model-free offline RL methods are widely studied to enhance the agent’s performance.
Model-based Offline RL. Model-based offline RL generates more state-action pair data through the trained environment model and improves the exploration for the OOD region. Since the gap between the trained and the actual environment model, the current model-based offline RL algorithms adopt a penalty, regularization or adversarial training method to perform a conservative policy. Yu et al. [26], Chen et al. [31] and Zhu et al. [38] estimated model uncertainty in different ways, and used model uncertainty as a reward penalty to punish model data. Yu et al. [28] considered the estimation errors for model uncertainty and utilized value function regularization to push down unreliable model data. Different from the above value penalty methods, Rigter et al. [36] and Bhardwaj et al. [37] used an adversarial training framework, forcing the agent to act conservatively in the OOD region. Shi et al. [39] and Liu et al. [40] incorporated a robust Markov Decision Process into policy learning to achieve conservative value estimation for model data.
Our work. The proposed method, DOMAIN, belongs to model-based offline RL. The current model-based offline RL methods seldom consider the difference between the model data, that is, the precise model data is rarely fully utilized, leading to overly conservative policies. Compared with previous works [26, 27, 28, 31, 36, 37], the DOMAIN adopts the adaptive sampling distribution of model data, where the penalty for high-error model data is increased and the penalty for accurate model data is reduced or even rewarded. Moreover, although training the discriminator to calculate the adaptive sampling distribution requires additional computational cost, the computational complexity of DOMAIN is lower than that of the adversarial training method [36, 37]. Theoretical analysis shows that DOMAIN has a policy improvement guarantee, and $Q$-values learned by DOMAIN for the OOD region are lower bounds of the true $Q$-values.
III Preliminaries
In this section, we provide the basic notations for reinforcement learning algorithms, including offline RL and model-based offline RL.
@@ III Preliminaries
Reinforcement Learning. The framework of RL is based on the Markov Decision Process (MDP) $\mathcal{M}=\left(\mathcal{S},\mathcal{A},r,T_{\mathcal{M}},\rho_{0},\gamma\right)$, where $\mathcal{S}$ is the set of states, $\mathcal{A}$ is the set of actions, $r(s,a)$ is the reward function, $T_{\mathcal{M}}\left(s^{\prime}\mid s,a\right)$ represents the dynamic system under $\mathcal{M}$, $\rho_{0}$ is the initial state distribution, and $\gamma\in(0,1)$ is the discount factor [10]. The goal of RL is to learn a policy that maximizes the cumulative reward $J(\mathcal{M},\pi)=\mathbb{E}_{(s,a)\sim d_{\mathcal{M}}^{\pi}(s,a)}[r(s,a)]/1-\gamma$, where $d_{\mathcal{M}}^{\pi}(s,a)=d_{\mathcal{M}}^{\pi}(s)\pi(a\mid s)$ is the marginal distribution of states and actions under the learned policy $\pi$, and $d_{\mathcal{M}}^{\pi}(s)$ is the discounted marginal state distribution under the learned policy $\pi$.
@@
Offline RL. The goal of offline RL is to learn a policy $\pi_{\theta}(\cdot\mid s)$ from an offline dataset $\mathcal{D}=\left\{\left(s_{i},a_{i},r_{i},s_{i}^{\prime}\right)\right\}_{i=1}^{n}$ without interacting with the environment, to maximize the expected discounted reward. However, there exists the distribution shift between the learned policy $\pi_{\theta}(\cdot\mid s)$ and the behavioral policy $\pi_{b}(\cdot\mid s)$ of the dataset, leading to unstable training of the agent [10]. To alleviate this issue, previous studies [17, 18, 19, 20, 21] have introduced regularization on the value function of OOD actions to reduce the probability of selecting OOD actions. Here primarily focuses on CQL[20]. The formulas for policy improvement and policy evaluation are as follows:
@@
Model-based Offline RL. Model-based offline RL involves two steps: environment model learning and policy learning. It uses the learned environment model $\hat{T}$ to generate more model data, which can broaden data coverage and can further enhance the exploration for the OOD region. The environment model $\widehat{T}_{\phi}\left(s^{\prime},r\mid s,a\right)$ is trained using the offline dataset, and then, the policy is learned based on both offline data and generated model data. Following the previous work [26, 27], the environment model is parameterized as Gaussian distribution $\widehat{T}_{\phi}(s^{\prime},r\mid s,a)=\mathcal{N}[\mu_{\phi}(s,a),\Sigma_{\phi}(s,a)]$, which can better describe the randomness of the real environment. The loss function for training the environment model employs maximum likelihood estimation, which can be denoted as: $\mathcal{L}_{\mathcal{M}}(\phi)=-\mathbb{E}_{\left(s,a,r,s^{\prime}\right)\sim\mathcal{D}}[\log\widehat{T}_{\phi}(s^{\prime},r|s,a)]$. With this learned model, an estimated MDP $\widehat{\mathcal{M}}$ can be obtained.
Subsequently, the agent interacts with the learned policy $\widehat{\pi}$ and the environment model $\widehat{T}_{\phi}$ to generate model data $\mathcal{D}_{\text{model }}$ through $H$-step rollouts. The policy learning is then performed using the combined dataset $\mathcal{D}_{\text{offline }}\cup\mathcal{D}_{\text{model }}$, where the data is sampled from model data $\mathcal{D}_{\text{model }}$ with the probability $f,f\in[0,1]$, and from the offline data $\mathcal{D}_{\text{offline }}$ with the probability $1-f$. However, offline dataset fails to cover the entire state-action space, resulting in disparities between the learned environment model $\widehat{T}_{\phi}\left(s^{\prime},r\mid s,a\right)$ and the actual environment model $T\left(s^{\prime},r\mid s,a\right)$. Similar to [35, 36], we employed a model ensemble approach to reduce biases introduced by probabilistic networks.
IV Methodology
This section provides a detailed description of the mildly conservative model-based offline reinforcement learning (DOMAIN) algorithm, including the mildly conservative estimation of the value function and adaptive value regularization on model data.
IV-A Mildly Conservative Value Estimation
Conservative policy evaluation within the DOMAIN algorithm framework can be described as follows: Given an offline dataset $\mathcal{D}_{\text{offline }}=\left\{\left(s_{i},a_{i},r_{i},s_{i}^{\prime}\right)\right\}_{i=1}^{n}$, a learned policy $\widehat{\pi}$, and a learned environmental model $\widehat{T}_{\phi}$, the primary objective is to obtain a conservative estimate of $Q(s,a)$. To achieve this, the policy evaluation step in DOMAIN utilizes the CQL framework (see Eq. (2)). This principle can be mathematically expressed as
@@ IV-A Mildly Conservative Value Estimation
where $f$ can be determined by two guidelines. One is the error of model data, and the other is the algorithm itself. If the model data is accurate and the algorithm can effectively penalize model data with large errors, $f$ can be set larger. On the contrary, if the error of model data is large and the algorithm fails to effectively penalize them, $f$ needs to be set smaller to guarantee the stability and security of the agent. $\widehat{\mathcal{T}}^{\pi}$ represents the actual Bellman operator used in the algorithm, while $\mathcal{T}_{\overline{\mathcal{M}}}^{\pi}$ and $\mathcal{T}_{\widehat{\mathcal{M}}}^{\pi}$ represent the operator applied under the empirical MDP and the learned model, respectively.
The value function $Q$ and the policy $\pi$ in Eq. (4) can be approximated using neural networks, denoted as $Q_{\omega}$ and $\pi_{\theta}$, respectively. The value function update follows the SAC framework [42], and consequently, the Bellman operator can be further expressed as:
@@
where $\alpha$ is the entropy regularization coefficient, which determines the relative importance of entropy. By employing a mildly conservative policy evaluation, the policy improvement scheme can be derived:
@@
where $Z$ is the normalization factor. Substituting the solution obtained by solving Eq. (7) back into Eq. (3), the following expression can be derived:
@@
IV-B Adaptive Value Regularization on Model Data
When performing policy evaluation, we aim to assign higher penalty coefficients to larger model data errors, thereby increasing the value for $\omega(s,a)$. Drawing inspiration from [43, 44], the KL divergence is employed to quantify the magnitude of model data errors: